find three consecutive odd integers such that the sum of the second and third integer is four times the first integer.
Answer by checkley77(12569)

Let x, x+2 & x=4 be the 3 integers.
x+2+x+4=4x
x+x-4x--4-2
-2x=-6
x=-6/-2
x=3 ans.for the smallest integer.
3+2=5 ans. for thr middle integer.
3+4=7 ans. for the larfest integer.
Proof:
5+7=4*3
12=12